# UNet with Axial Transformer : A Neural Weather Model for Precipitation Nowcasting

### v1

@@ Abstract
Making accurate weather predictions can be particularly challenging for localized storms or events that evolve on hourly timescales, such as thunderstorms. Hence, our goal for the project was to model Weather Nowcasting for making highly localized and accurate predictions that apply to the immediate future replacing the current numerical weather models and data assimilation systems with Deep Learning approaches. A significant advantage of machine learning is that inference is computationally cheap given an already-trained model, allowing forecasts that are nearly instantaneous and in the native high resolution of the input data. In this work we developed a novel method that employs Transformer-based machine learning models to forecast precipitation. This approach works by leveraging axial attention mechanisms to learn complex patterns and dynamics from time series frames. Moreover, it is a generic framework and can be applied to univariate and multivariate time series data, as well as time series embeddings data. This paper represents an initial research on the dataset used in the domain of next frame prediciton, and hence, we demonstrate state-of-the-art results in terms of metrices (PSNR = 47.67, SSIM = 0.9943) used for the given dataset using UNet with Axial Transformer. The code is available at https://github.com/sumitmamtani/Precipitation-Nowcasting
@@ 1 Introduction
Deep Neural Networks (DNN)[17, 13, 1, 12] have been successfully applied in many diverse domains such as image classification, video analysis, language modeling and translation, medical imaging, and weather. Recently, there has been increasing interest in using DNNs to generate and improve weather forecasting which is an unsupervised representation problem. In these kinds of problems, next-frame prediction is a new, promising direction of research in computer vision, predicting possible future images by presenting historical image information.
@@
Weather forecasting is the prediction of future weather conditions such as precipitation, temperature, pressure, and wind and is fundamental to both science and society. Our particular interest is in the area of nowcasting, a term used to describe high resolution, short-term (e.g. 0 to 2 hours), weather forecasts of precipitation, or other meteorological quantities. The precipitation nowcasting field helps in the accurate prediction of rainfall over an area by looking at radar images. The field deals with the generation of the radar image at some points in the near future.
@@
One of the various architectures researchers have tried to implement is Conv-LSTM[14, 16, 8] that has shown great results in dealing with time-series data due to them being pretty good at extracting patterns in input feature space, where the input data spans over long sequences. Given the gated architecture of LSTM’s that has this ability to manipulate their memory state, they are ideal for such problems.
@@
With the continuous input of data from one end, we know that the prediction needs to be swift while dealing with a huge amount of data. With the same thing in mind, some recent research introduced UNet to counter the same issue and were successful to get good results on image to image translation problems[2, 19].
@@
Usage of conditional GANs (cGANs)[6] has became widely popular under the domain of image-to-image translation. As we require a model that would learn from previous inputs in the sequence, we tried to make use of this property where cGANs predicts N future radar frames given M past-conditional frames.
@@
In this Paper, we also introduced encode-decoder architecture[7] with the Axial Transformer[4] in our Novel model, a simple yet effective self-attention-based [20, 21] autoregressive model for data organized as multidimensional tensors. Rather than applying attention to a flattened string of tensor elements, Axial Transformer instead applies attention along a single axis of the tensor without flattening, so this is referred to as “axial attention.” Since the length of any single axis (that is, the height or width of an image) is typically much smaller than the total number of elements, an axial attention operation enjoys a significant saving in computation and memory over standard self-attention. We achieve state-of-the-art results by training the encode-decoder with Axial Transformer framework. Our contributions are summarized as follows:
  • We implemented an encoder decoder architecture and passed the frames of videos as input to the architecture.
  • The output of encoder decoder architecture is passed as input to the Axial Transformer network.
  • With sequence frames given as input, the final output of the UNet with Axial Transformer model is a distribution, out of which we select the mean value for every pixel.
@@ 2 Related Work
Previously, the Conv-LSTM model has shown promising results on the next frame prediction problem on the Moving MNIST dataset [14]. The dataset was initially created in the context of Unsupervised Learning of Video Representations [15] and used LSTM to learn the representation of video sequence. As the radar images too are a kind of time series data and we know that clouds can not abruptly change direction or disappear, we know that there are some motion parameters associated with a continuous sequence of radar images too. Using the same idea of next frame prediction in a video sequence, ConvLSTM can be applied to a dataset of radar images too.
UNet is one of the recent models that has showcased its ability to predict the next frame of time-series data very well [2]. In the paper, as part of Traffic4cast challenge 2019, UNet was used to predict short-term traffic flow volume. The input and output of the model were the same sizes that are also relevant in our context as we need to reproduce radar images of the same location sometime in the future. Another paper that actually implemented UNet for the weather forecasting problem [19] saw a significant increase in results on the dataset consisting of precipitation maps from a region of the Netherlands and a binary image of cloud coverage of France. The size of the model here is very small compared to previous models that were used to solve the same problem, which is also a significant advantage considering the latency requirement of our problem statement.
Researchers have now tried an observations-driven approach for probabilistic nowcasting using deep generative models (DGMs). DGMs are statistical models that learn probability distributions of data and allow for easy generation of samples from their learned distributions[11, 9]. As generative models are fundamentally probabilistic, they have the ability to simulate many samples from the conditional distribution of future radar given historical radar.
@@
3 Our Method
We next describe the neural network architectures used in the expirement of this project in the following section:
3.1 Conv-LSTM
This model was our baseline model which was inspired by the paper that implemented the same model on the original complete HKO-7 dataset, along with precipitation data in .csv and .pkl files [14]. The PyTorch implementation of the code is available on GitHub  [^1] https://github.com/sladewinter/ConvLSTM. As the code mentioned worked on the Moving MNIST dataset, we had to integrate our dataset running for the given model. The model architecture was suitable to use in the context of this problem, as the original paper too works on the next frame prediction problem and learns the representation of video sequence. By extending the fully connected LSTM (FC-LSTM) to have convolutional structures in both the input-to-state and state-to-state transitions, the paper proposes the convolutional LSTM (ConvLSTM) and used it to build an end-to-end trainable model for the precipitation nowcasting problem. The sequence to sequence model acts as an encoder-decoder structure with 64 kernels each with size of 3 and padding of 1, to ensure the input and output of the model stay of equal size. The activation function used was ReLU and optimizer being Adam.
Each of the 3 sequential layers used was ConvLSTM cell which consisted of a Conv-2D layer, input, output, and forget gate all having a Sigmoid activation in the LSTM cell. Each of the gates takes the hidden state and the current input as input, it concatenates the vectors and applies the sigmoid activation, now the input gate controls if the memory cell should be updated. Forget gate controls how much of the old state should be forgotten. The LSTMs are able to learn sequential data due to the above properties.
3.2 Conditional generative adversarial network (cGANs)
@@
GANs are generative models that learn a mapping from random noise vector z to output image y, G : z → y. But conditional GANs learn a mapping from observed image x and random noise vector z, to y, G : x, z → y. The generator G is trained such that it produces outputs that cannot be distinguished from “real” images by a trained discriminator (D) which is trained to do as well as possible at detecting the generator’s fakes images.
@@
where G tries to minimize this objective against an adversarial D that tries to maximize it, so overall objective function is given by:-
@@
$L(G,D)$ is known as adversarial loss. We have also used Pixelwise L1 content loss to train generator and discriminator model. So the generator task is to not only fool the discriminator but also to be near the ground truth output in an L1 sense. So our total loss comprises of addition of both the pixelwise loss and adversarial loss. We have used L1 distance rather than L2 as L1 encourages less blurring. So L1 loss are as follows is given by :
$\begin{split}L_{L1}(G)\ =\ E_{x,y}[\ ||\ y\ -\ G(x,z)\ ||_{1}]\end{split}$ (3)
@@
where $\lambda$ is a hyper-parameter and in our implementation we have set its value to 100.
@@ 3.2.1 Generator
The generator consists of a series of layers that progressively downsample until a bottleneck layer is reached, after that we reverse the same process i.e we upsample. Such a network requires that all information flow passes through all the layers, including the bottleneck. So there is a low level of information shared like the prominent edge locations between input and output which produces high-quality images than simple CNN generator model.
We have added skip connections, to avoid the vanishing gradient problem and to incorporate information from earlier layers. This is the general shape of a UNet architecture. Specifically, we add skip connections between each layer $i$ and layer $n-i$, where $n$ is the total number of layers. Each skip connection simply concatenates all channels at layer $i$ with those at layer $n-i$.
@@ 3.2.2 Discriminator
The discriminator is a type of Convolution Neural Network(CNN) model. The discriminator has 4 sequential blocks having 64, 128, 256 , 512 filters respectively. Each sequential block consists of Convolutional layer, Batch-normalization layer and activation layer as LeakyRelu. The generator tries to fool the discriminator by creating fake images. The discriminator is trained in such a way that it tries to distinguish real data from the data created by the generator.
Figure 1: The figure shows the final novel UNet with Axial Transformer model. Part (1) is our UNet model that encompasses encoder and decoder using Conv2D. Part (2) is the transformer [18] that attends every row and column for each location (pixel), hence an axial attention [4]. Part (3) is the distribution we get as our final output of which we take mean to make final prediction for Tth minute
3.3 UNet
UNets, first introduced in Convolutional Networks for Biomedical Image Segmentation paper[10], have been able to expand their use case from image segmentation to predicting the future sequence too. The architecture consists of a contracting path to capture context and a symmetric expanding path that enables precise localization. Encoder (downsampling path) extracts a meaningful feature map from an input image. As is standard practice for a CNN, the Encoder, doubles the number of channels at every step and halves the spatial dimension. Next, the Decoder (upsampling path) actually upsamples the feature maps, where at every step, it doubles the spatial dimension and halves the number of channels (opposite to that of what an Encoder does).
The contractive path (Encoder) consists of the repeated application of two 3x3 convolutions (padding=1), each followed by a rectified linear unit (ReLU) and a 2x2 max pooling operation with stride 2 for downsampling. At each downsampling step we double the number of feature channels. As the image size we use (128 x 128) is very small to be downsampled too much, we couldnt traverse the contractive path as much as given in the reference papers, i.e. rather than having out channels as 1024 in final output of our Encoder, we had 256 as our number of out channels. Hence, the sequence of in and out channels for each block in the Encoder consisted of [2,64,128,256].
The decoder layer is explained as following: every step in the expansive path consists of an upsampling of the feature map followed by a 2x2 convolution (up-convolution) that halves the number of feature channels, a concatenation with the correspondingly cropped feature map from the contracting path, and two 3x3 convolutions, each followed by a ReLU. The cropping is necessary due to the loss of border pixels in every convolution. The ConvTranspose2d operation performs the up-convolution and again the same block consisting of Conv2D and ReLU in between is used to half the number of channels. The output of this Decoder is 128 x 128 image for each batch, which can be again compared to the target and initiate the learning of our model.
The usage of residual skip connections helps alleviate the vanishing gradient problem allowing for UNet models with deeper neural networks to be designed. Each residual unit can be denoted by the following expressions:
$y_{l}=h(x_{l})\ +F(x_{l},w_{l})x_{l+1}=y_{l}$ (5)
$x_{l+1}=y_{l}$ (6)
To handle a time series data that we have, we could have probably used Conv-3D and MaxPool-3D layers in our UNet. But as the data was preprocessed into a GRAYSCALE image, we could just substitute the number of frames as the number of input channels, an idea inspired by SmaAt-UNet. [19].
3.4 UNet with Axial Transformer
Our Novel model is based on axial attention, a simple generalization of self-attention that naturally aligns with the multiple dimensions of the tensors in both the encoding and the decoding settings. As we know Attention mechanisms have become an integral part of compelling sequence modeling and transduction models in various tasks, allowing modeling of dependencies without regard to their distance in the input or output sequences. So we have used Axial Transformers, an axial-attention-based autoregressive model for images and other data organized as high dimensional tensors.
The sequence of images, which are the first 16 frames from the sequence of length 20, are fed into our Downsampler, which is same as the Encoder of our UNet model used before. This acts as an image processing pipeline and the first stage of our model. The next layer is a Decoder layer that outputs the a representation of the time series, i.e. integrating the information over-time. The Internal layers of both Encoder and Decoder are made of several layers of CNN. We produce one feature map per input image from the sequence of length 16, previously in the simple UNet model we compressed all information within a single image output from the model, but now these 16 images will act as an input to our next stage- Axial Transformer [4].
Axial attention can be used within standard Transformer layers in a straightforward manner to produce Axial Transformer layers. The basic building blocks are the same as those found in the standard Transformer architecture:
$h\leftarrow\text{Embed}(x)$
[…]
$h\leftarrow\text{MaskTransformerBlock2}(h)\times L_{\text{row}}$
$L_{row}$ is the number of masked row attention blocks applied to h, The operation ShiftRight shifts the input right by one pixel. PositionEmbeddings is a tensor of position embeddings that inform the attention layers of the position. x is the gray scale frame.
Outer Decoder: Each pixel in the model depends on previous pixels in its own row. To capture all previous rows, insert unmasked row and masked column layers in the beginning of the model as described below:
$h\leftarrow\text{Embed}(x)$
[…]
$u\leftarrow\text{MaskTransformerBlock1}(\text{Block2}(u))\times\frac{L_{\text{upper}}}{2}$
[…]
$h\leftarrow\text{MaskedTransformerBlock2}(h)\times L_{\text{row}}$
The idea of using a transformer with a Encoder-Decoder model was inspired by a recent development in the field of Precipitation Nowcasting at Google who introduced a Neural Weather model called MetNet [18].
Now what Axial Tranformer will try to do is, encode information in these feature maps from the space around each point of interest. Hence, attending parts that are relevant to justify motion of a cloud patch in a radar image. The way Transformer works is- we have a series of feature maps that are continuous with respect to time. Each pixel from the latest frame emits a query vector and each of the pixel from the older image feature maps emits a key. And each of the pixel emitting query can look at (attend) each of the pixel in the lower layer (keys). Hence, we are able to incorporate long range dependencies aggregating information from the downstream and increasing the resolution. But in Axial Transformers we will specifically attend the pixels that are either in the same row or same column in the images produced. This saves the required memory for the computation over the layers of attention while aggregating information over the spatial dimensions. The code implementation of the Transformer was inspired by the PyTorch implementation of axial attention available on GitHub  [^2] https://github.com/lucidrains/axial-attention.
The final output of the this model is a distribution across 128 frames output from transformer head out of which we need to take the mean value as the predicted state of a pixel after the prediction. This predicted tensor, flattened, along with the target image tensor also flattened, is sent to calculate the loss and start model learning.

### v2

# Axial-UNet: A Neural Weather Model for Precipitation Nowcasting

@@ Abstract
Accurately predicting short-term precipitation is critical for weather-sensitive applications such as disaster management, aviation, and urban planning. Traditional numerical weather prediction can be computationally intensive at high resolution and short lead times. In this work, we propose a lightweight UNet-based encoder–decoder augmented with axial-attention blocks that attend along image rows and columns to capture long-range spatial interactions, while temporal context is provided by conditioning on multiple past radar frames. Our hybrid architecture captures both local and long-range spatio-temporal dependencies from radar image sequences, enabling fixed lead-time precipitation nowcasting with modest compute. Experimental results on a preprocessed subset of the HKO-7 radar dataset demonstrate that our model outperforms ConvLSTM, pix2pix-style cGANs, and a plain UNet in pixel-fidelity metrics, reaching PSNR 47.67 and SSIM 0.9943. We report PSNR/SSIM here; extending evaluation to meteorology-oriented skill measures (e.g., CSI/FSS) is left to future work. The approach is simple, scalable, and effective for resource-constrained, real-time forecasting scenarios.
@@ 1 Introduction
Deep neural networks (DNN)[21, 17, 2, 16, 12] have been successfully applied in many diverse domains such as image classification, video analysis, language modeling and translation, medical imaging, and weather. Recently, there has been increasing interest in using DNNs to generate and improve weather forecasting, which is an unsupervised representation problem. In these kinds of problems, next-frame prediction is a new, promising direction of research in computer vision, predicting possible future images by presenting historical image information. Recent neural network-based weather models highlight this trend toward learned forecasting [3].
@@
Weather forecasting is the prediction of future weather conditions, such as precipitation, temperature, pressure, and wind, and is fundamental to both science and society. Our particular interest is in the area of nowcasting, a term used to describe high-resolution, short-term (e.g., 0 to 2 hours) weather forecasts of precipitation or other meteorological quantities. The precipitation nowcasting field helps in the accurate prediction of rainfall over an area by looking at radar images. The field deals with the generation of the radar image at some points in the near future.
@@
One of the various architectures researchers have tried to implement is Conv-LSTM[18, 20, 11], which has shown great results in dealing with time-series data because it is pretty good at extracting patterns in the input feature space, where the input data spans over long sequences. The gated architecture of LSTMs has the ability to manipulate the memory state, making it more subtle for such problems.
@@
With the continuous input of data from one end, we know that the prediction needs to be swift while dealing with a huge amount of data. Following this, some recent works have used UNet for this problem and shown improved results for image-to-image translation problems [4, 23].
@@
Recently, conditional GANs (cGANs) [9] have become widely popular under the domain of image-to-image translation[7]. As we require a model that would learn from previous inputs in the sequence, we tried to make use of this property, where cGANs predict N future radar frames given M past-conditional frames.
@@
In this Paper, we also introduced an encoder-decoder architecture [10] with the Axial Transformer [6] in our proposed model, a simple yet effective self-attention-based [24, 26] autoregressive model for data organized as multidimensional tensors. Rather than applying attention to a flattened string of tensor elements, Axial Transformer instead applies attention along a single axis of the tensor without flattening, so this is referred to as “axial attention”. Since the length of any single axis (that is, the height or width of an image) is typically much smaller than the total number of elements, an axial attention operation enjoys a significant saving in computation and memory over standard self-attention. We achieve competitive results on a preprocessed dataset by training the encoder-decoder with the Axial Attention blocks. Our contributions are summarized as follows:
  • Efficient encoder–decoder for fixed lead-time nowcasting. We propose a lightweight encoder–decoder that ingests $M$ past radar frames and predicts the next frame at a fixed lead time, maintaining native resolution while keeping compute modest.
  • Axial attention for long-range spatial context. We insert axial-attention blocks that attend along rows and columns, capturing large-scale advection patterns with near-linear memory; temporal context is supplied by the multi-frame encoder.
  • Controlled evaluation on an HKO-7 subset. All models share the same preprocessing (grayscale $128\times 128$) and the same train/val/test split. Each baseline uses its typical number of input frames—ConvLSTM (15), cGAN (4), and UNet/ours (16)—and we evaluate fixed lead-time (next-frame; $M\!\to\!1$) predictions, extending to longer horizons via autoregression.
  • Quality gains. The axial-attention variant improves PSNR/SSIM over strong baselines(see Table 2).
@@ 2 Related Work
Previously, the Conv-LSTM model has shown promising results on the next frame prediction problem on the Moving MNIST dataset [18]. The dataset was initially created in the context of Unsupervised Learning of Video Representations [19] and used LSTM to learn the representation of the video sequence. As the radar images are also a kind of time series data, and we know that clouds can not abruptly change direction or disappear, we know that there are some motion parameters associated with a continuous sequence of radar images, too. Using the same idea of next frame prediction in a video sequence, ConvLSTM can also be applied to a dataset of radar images.
UNet is one of the recent models that has shown its ability to predict the next frame of time-series data very well [4]. In the paper, as part of the Traffic4cast challenge 2019, UNet was used to predict short-term traffic flow volume. The input and output of the model were the same sizes, that are also relevant in our context, as we need to reproduce radar images of the same location sometime in the future. Another paper that actually implemented UNet for the weather forecasting problem [23] saw a significant increase in results on the dataset consisting of precipitation maps from a region of the Netherlands and a binary image of cloud coverage of France. The size of the model here is very small compared to previous models that were used to solve the same problem, which is also a significant advantage, considering the latency requirement of our problem statement.
Researchers have now tried an observations-driven approach for probabilistic nowcasting using deep generative models (DGMs). DGMs are statistical models that learn probability distributions of data and allow for easy generation of samples from their learned distributions[15, 12]. As generative models are fundamentally probabilistic, they can simulate many samples from the conditional distribution of future radar given historical radar.
@@
Summary and gap.
ConvLSTM- and GAN-based approaches capture temporal dynamics but struggle with long-range spatial dependencies or require heavy memory. Plain UNet models are efficient but lack explicit mechanisms for large-scale advection. This motivates our design: a lightweight encoder–decoder augmented with axial attention to model row/column interactions efficiently.
3 Method: UNet with Axial Attention (Axial-UNet)
Given $M$ past radar frames (here M=16M{=}16) sampled every 6 minutes, our model predicts the next frame at a fixed lead time; longer horizons are obtained via autoregression. The backbone is a UNet-style encoder–decoder, and we insert axial-attention blocks that attend along rows and columns to capture large-scale advection with modest memory overhead.
3.1 Backbone encoder–decoder
UNets, first introduced in the Convolutional Networks for Biomedical Image Segmentation paper[14], have been able to expand their use case from image segmentation to predicting the future sequence too. The architecture consists of a contracting path to capture context and a symmetric expanding path that enables precise localization. Encoder (downsampling path) extracts a meaningful feature map from an input image. As is standard practice for a CNN, the Encoder doubles the number of channels at every step and halves the spatial dimension. Next, the Decoder (upsampling path) actually upsamples the feature maps, where at every step, it doubles the spatial dimension and halves the number of channels (opposite to what an Encoder does).
The contractive path (Encoder) consists of the repeated application of two 3x3 convolutions (padding=1), each followed by a rectified linear unit (ReLU) and a 2x2 max pooling operation with stride 2 for downsampling. At each downsampling step, we double the number of feature channels. As the image size we use (128 x 128) is very small to be downsampled too much, we couldn’t traverse the contractive path as much as given in the reference papers, i.e., rather than having out channels as 1024 in the final output of our Encoder, we had 256 as our number of out channels. Hence, the sequence of in and out channels for each block in the Encoder consisted of [16,64,128,256].
The decoder layer is explained as follows: every step in the expansive path consists of an upsampling of the feature map followed by a 2x2 convolution (up-convolution) that halves the number of feature channels, a concatenation with the correspondingly cropped feature map from the contracting path, and two 3x3 convolutions, each followed by a ReLU. The cropping is necessary due to the loss of border pixels in every convolution. The ConvTranspose2d operation performs the up-convolution, and again the same block consisting of Conv2D and ReLU in between is used to half the number of channels. The output of this Decoder is a 128 x 128 image for each batch, which can again be compared to the target and initiate the learning of our model.
The usage of residual skip connections helps alleviate the vanishing gradient problem, allowing for UNet models with deeper neural networks to be designed. Each residual unit can be denoted by the following expressions:
$y_{\ell}=h(x_{\ell})+F(x_{\ell},W_{\ell}),\quad x_{\ell+1}=y_{\ell}.$ (1)
To handle the time series data that we have, we could have probably used Conv-3D and MaxPool-3D layers in our UNet. But as the data was preprocessed into a GRAYSCALE image, we could just substitute the number of frames as the number of input channels, an idea inspired by SmaAt-UNet. [23]. Recent UNet variants with attention also show benefits in remote sensing nowcasting [27].
3.2 Axial attention blocks
Our proposed model is based on axial attention, a simple generalization of self-attention that naturally aligns with the multiple dimensions of the tensors in both the encoding and the decoding settings. As we know, Attention mechanisms have become an integral part of compelling sequence modeling and transduction models in various tasks, allowing modeling of dependencies without regard to their distance in the input or output sequences. So we have used Axial Transformers, an axial-attention-based autoregressive model for images and other data organized as high-dimensional tensors.
The sequence of images, which are the first 16 frames from the sequence of length 20, is fed into our Downsampler, which is the same as the Encoder of our UNet model used before. This acts as an image processing pipeline and the first stage of our model. The next layer is a Decoder layer that outputs a representation of the time series, i.e., integrating the information over time. The Internal layers of both Encoder and Decoder are made of several layers of CNN. We produce one feature map per input image from a sequence of length 16. Previously, in the simple UNet model, we compressed all information within a single image output from the model, but now these 16 images will act as input to our next stage, the Axial Transformer [6]. We adopt axial attention to model long-range spatial interactions with reduced memory cost in radar imagery [25].
Axial attention can be straightforwardly used within standard Transformer layers to produce Axial Transformer layers. The theoretical foundation for axial attention is established in prior work [6], which provides the mathematical formulation and proof for decomposing two-dimensional self-attention into separable row and column-wise operations. In this study, we adopt their proven formulation within our encoder-decoder framework, demonstrating its empirical stability and effectiveness for spatio-temporal precipitation nowcasting. The basic building blocks are the same as those found in the standard Transformer architecture:
h ← Embed(x)
h ← ShiftRight(h) + PositionEmbeddings
$h\leftarrow\text{MaskTransformerBlock2}(h)\times L_{\text{row}}$
$L_{row}$ is the number of masked row attention blocks applied to h. The operation ShiftRight shifts the input right by one pixel. PositionEmbeddings is a tensor of position embeddings that inform the attention layers of the position. x is the gray scale frame.
Outer Decoder: Each pixel in the model depends on previous pixels in its own row. To capture all previous rows, we insert unmasked row and masked column layers at the beginning of the model as described below:
h ← Embed(x)
u ← h + PositionEmbeddings
u ← MaskTransformerBlock1(Block2(u)) × $L_{upper}/2$
h ← ShiftDown(u) + ShiftRight(h) + PositionEmbeddings
$h\leftarrow\text{MaskedTransformerBlock2}(h)\times L_{\text{row}}$
The idea of using a transformer with an encoder-decoder model was inspired by a development in the field of Precipitation Nowcasting at Google, which introduced a Neural Weather model called MetNet [22].
Now, what Axial Transformer will try to do is encode information in these feature maps from the space around each point of interest. Hence, attending to parts that are relevant to justify the motion of a cloud patch in a radar image. The way the Transformer works is - we have a series of feature maps that are continuous with respect to time. Each pixel from the latest frame emits a query vector, and each of the pixels from the older image feature maps emits a key. And each of the pixel-emitting queries can look at (attend) each of the pixels in the lower layer (keys). Hence, we can incorporate long-range dependencies by aggregating information from the downstream and increasing the resolution. But in Axial Transformers, we will specifically attend to the pixels that are either in the same row or the same column in the images produced. This saves the required memory for the computation over the layers of attention while aggregating information over the spatial dimensions. The code implementation of the Transformer was inspired by the PyTorch implementation of axial attention available on GitHub  [^1] https://github.com/lucidrains/axial-attention.
The final output of this model is a distribution across 128 frames, output from the transformer head, out of which we need to take the mean value as the predicted state of a pixel after the prediction. This predicted tensor, flattened, along with the target image tensor, also flattened, is sent to calculate the loss and start model learning.
3.3 Baselines
We compare against (i) ConvLSTM and (ii) a pix2pix-style conditional GAN (cGAN), in addition to a plain UNet backbone. All baselines share our preprocessing and split (grayscale $128{\times}128$). Input lengths follow customary practice—ConvLSTM (15), cGAN (4), UNet/ours (16)—and we evaluate fixed lead-time ($M\!\to\!1$) predictions, extending to multiple frames via autoregression.
3.3.1 ConvLSTM
We use ConvLSTM as a baseline, following Shi et al. [18] on the HKO-7 dataset. We adapt a PyTorch implementation [^2] https://github.com/sladewinter/ConvLSTM originally developed for Moving MNIST to our radar sequences. ConvLSTM extends fully connected LSTMs by replacing affine transforms with convolutions in both the input-to-state and state-to-state transitions, enabling end-to-end sequence-to-sequence modeling for precipitation nowcasting.
Our implementation uses a three-layer encoder–decoder with 64 $3\times 3$ kernels per layer and padding 1 to preserve spatial size, ReLU activations, and the Adam optimizer. Each ConvLSTM cell comprises a Conv2D layer and input, output, and forget gates with sigmoid activations; the gates control memory updates and retention of past state. In our setup we condition on M=15M{=}15 input frames and predict the next frame, rolling out longer horizons autoregressively. For a recent survey of deep learning for precipitation nowcasting and datasets/metrics, see [1].
3.3.2 cGANs
@@ 3.3.1 ConvLSTM
GANs are generative models that learn a mapping from a random noise vector z to an output image y, G: z → y. But conditional GANs learn a mapping from observed image x and random noise vector z, to y, G: x, z → y. The generator G is trained such that it produces outputs that cannot be distinguished from “real” images by a trained discriminator (D), which is trained to do as well as possible at detecting the generator’s fake images.
@@
where G tries to minimize this objective against an adversarial D that tries to maximize it, so the overall objective function is given by:-
@@
Figure 1: Architecture of Axial-UNet (UNet with axial attention). Part (1) is our UNet model that encompasses an encoder and decoder using Conv2D. Part (2) is the transformer [22] that attends every row and column for each location (pixel), hence an axial attention [6]. Part (3) is the distribution we get as our final output, of which we take the mean to make a final prediction for Tth minute
$L(G,D)$ is known as adversarial loss. We have also used Pixelwise L1 content loss to train the generator and discriminator model. So the generator task is to not only fool the discriminator but also to be near the ground truth output in an L1 sense. So our total loss comprises the addition of both the pixelwise loss and adversarial loss. We have used L1 distance rather than L2 as L1 encourages less blurring. So L1 loss are as follows is given by:
[…]
@@
where $\lambda$ is a hyper-parameter, and in our implementation, we have set its value to 100.
@@ 3.3.3 Generator
The generator consists of a series of layers that progressively downsample until a bottleneck layer is reached, after that we reverse the same process, i.e, we upsample. Such a network requires that all information flow passes through all the layers, including the bottleneck. So there is a low level of information shared, like the prominent edge locations between input and output, which produces higher-quality images than a simple CNN generator model.
We have added skip connections to avoid the vanishing gradient problem and to incorporate information from earlier layers. This is the general shape of a UNet architecture. Specifically, we add skip connections between each layer $i$ and layer nin-i, where $n$ is the total number of layers. Each skip connection simply concatenates all channels at layer $i$ with those at layer nin-i.
@@ 3.3.4 Discriminator
The discriminator is a type of Convolution Neural Network(CNN) model. The discriminator has 4 sequential blocks having 64, 128, 256, 512 filters respectively. Each sequential block consists of the convolutional layer, batch-normalization layer, and activation layer as LeakyReLU. The generator tries to fool the discriminator by creating fake images. The discriminator is trained in such a way that it tries to distinguish real data from the data created by the generator.